# BODex: Scalable and Efficient Robotic Dexterous Grasp Synthesis Using Bilevel Optimization

### v1

@@ V-D2 Comparison of different dataset sizes
We explore the effect of dataset size on model performance, as shown in Fig. 7. The performance curve for datasets with objects on a table is more flat compared to floating objects, likely due to the more constrained distribution of the hand root pose — the hand is limited to positions over the table. The results also show that increasing the grasp number in the training dataset consistently leads to performance gains, though the trend slows down around 1 million. Future work could study enhancing model capacity to better leverage larger datasets.
@@ V-E Real-World Experiments
For each trial, we first segment the object on RGB by Segment Anything [38] and feed the segmented depth point cloud into the trained model. Next, a collision-free trajectory is planned and executed. Due to the noise in camera calibration, depth sensing, and the network output, we find that $1cm$ margin is not enough in the real world, so we instead use $\mathbf{x}_{p}^{\prime}=2\mathbf{x}_{p}-\mathbf{x}$ as the target for planning. Finally, the hand is controlled to the squeezed pose $\mathbf{x}_{s}$ and lifts the object.
@@
Figure 8: Real-world experiment setup. (Left) 20 test objects.
@@ VI Limitations and Future Work
First, our pipeline requires designating contact spheres on the hand, which we have placed on each fingertip. So the generated grasps rely solely on fingertips and do not utilize the palm. While fingertip-only grasps may facilitate future studies on tactile feedback, they lack the robustness that palm contact provides. Second, the collision-free trajectories in our dataset are not currently utilized, which may enable the training of a closed-loop visual policy in the future.

### v2

@@ V-D2 Comparison of different dataset sizes
We explore the effect of dataset size on model performance, as shown in Fig. 7. The performance curve for datasets with objects on a table is flatter compared to that for floating objects, likely due to the more constrained distribution of the hand root pose—restricted to positions over the table. The results also show that increasing the number of grasps in the training dataset consistently improves performance, though the gains slow down once the dataset scale reaches the million level. Future work could explore enhancing model capacity to better utilize larger datasets.
@@ V-E Real-World Experiments
For each trial, we first segment the object in the RGB image using Segment Anything [36] and feed the segmented depth point cloud into the trained model. The network then outputs 100 pairs of pre-grasp and grasp poses. Each pre-grasp pose is sequentially used as the target for collision-free motion planning with cuRobo, ordered by their probabilities. The first successfully planned pose is executed. Due to the noise in camera calibration, depth sensing, and network output, we find that a $1cm$ margin is insufficient in the real world, so we instead use $\mathbf{x}_{p}^{\prime}=2\mathbf{x}_{p}-\mathbf{x}$ as the new pre-grasp for motion planning. Finally, the hand moves to the squeezed pose $\mathbf{x}_{s}$ to apply force to the object before lifting it.
@@
Figure 8: Real-world experiment setup. Left: 20 test objects.
@@ VI Limitations and Future Work
First, our pipeline requires designating contact spheres on the hand, which we have placed on each fingertip. So the generated grasps rely solely on fingertips and do not utilize the palm. While fingertip-only grasps may facilitate future studies on tactile feedback, they lack the robustness that palm contact provides. Second, our work does not study grasping in a cluster scene. Third, our generated grasps are not functional. Finally, the collision-free trajectories in our dataset are not currently utilized, which may enable the training of a closed-loop visual policy in the future.